what is 2x+y=21 and x-y=6

Found 2 solutions by ikleyn, fractalier:
Answer by ikleyn(52778)

.
what is 2x+y=21 and x-y=6
------------------------------

It is the system of two equations in two unknowns.


Answer by fractalier(6550)

From these two equations, we add them to get
2x+y=21
x-y=6
-----------
3x = 27
x = 9
y = 3